solve by completing the square:
12x^2 + 17x=7





I got to the point :
x^2 + 17/12x + 289/256 = 7/12 + 286/256
but I dont understand how to simplify it

Answer by KnightOwlTutor(293)

12x^2 + 17x=7
Divide each term by 12 not 12x
x^2+17/12x=7/12
Then you half the second term and then square it then add this term to the other side
x^2 +17/24x +(17/24)^2=7/12 +(289/576)
Factor the left side into
(x+17/24)^2=(48+289/576=337/576
Take the square root of both sides REMEMBER THAT THE ANSWER CAN BE EITHER + OR -

x+17/24=+ or -(sqrt 337)/24
subtract 17/24 from both sides

x=-17-(sqrt 337)/24 0R -17+(sqrt 337)/24